OpenStudy (anonymous):

Factor. Check by multiplying factors 8x^3+2x^2+12x+3

OpenStudy (anonymous):

Set the equation to = 0. 8x^3 + 2x^2 + 12x + 3 = 0

OpenStudy (anonymous):

what? o_o

OpenStudy (anonymous):

Hero is about to explain. :D

hero (hero):

Factor by grouping: Notice that 8x^3 + 2x^2 = 2x^2(4x + 1) 12x + 3 = 3(4x + 1) That leaves something common to both factorizations.

OpenStudy (anonymous):

4x+1

hero (hero):

Can you figure out the complete factorization from here?

OpenStudy (anonymous):

2x^2(4x+1)= 8x^2+2x^2 3(4x+1)= 12x+3

hero (hero):

Start from this form: 8x^3 + 2x^2 + 12x + 3

hero (hero):

Your end result needs to be in the form (ax + b)(cx^2 + d)

OpenStudy (anonymous):

Where do I start first?

OpenStudy (anonymous):

@JungHyunRan

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@iGreen